OpenStudy (anonymous):

1. How many electrons reside on each pith ball?

OpenStudy (anonymous):

my class is online and my teacher is INCREDIBLY hard to understand. i'm not sure how to find this after doing the pith ball experiment....

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kainui):

Well @saiken2009 can you tell us what you know so far? Maybe draw a picture of the experiment and try to label things and try to relate the force pushing the balls apart to the charge of an electron.

OpenStudy (anonymous):

i got the equation Q=n*e, but not sure what each of those mean. what to plug in where. i got the pith ball = 6.168e^-9 coulomb after discharge and the measurments there. gotta answer these two questions : Questions: 1. How many electrons reside on each pith ball? 2. What is the gravitational attraction between the two pith balls? i think the gravitational attraction is f=(m1*m2*9.8)/r^2. but not too sure on that either. i've been trying to learn from khan academy videos.

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Ok, so it looks like you're on the right track so far, I'll help you understand this equation to begin with, Q=n*e. Q is the charge, which is in coulombs n is the number of electrons e is the charge of a single electron. So why does this equation work or matter? It's not so bad, this picture might help:  Let's pretend the ball has 5 electrons on it, so n=5. Each electron has a charge of 1.60*10^(-19) coulombs, so e=1.60*10^(-19) so if we add them all up, that means there are 5 of these electron charges, which is the total charge Q, so for my example this ends up being 8*10^(-19). Try to understand what each part of an equation means and I'll keep helping you to understand.

OpenStudy (anonymous):

ok, so my Q here is the 6.168e^-9 coulomb. and e i believe is always that 1.6e^-19 (constant i think he said....he's hard to understand). just solve for n?

OpenStudy (kainui):

One little thing to add to the last post: --- If we were to write out the units of the formula Q=n*e they would also work out, I will put brackets around the variable to mean "units of this thing" \[\Large [Q] = Coulombs \\ \Large [n] = number \ of \ electrons \\ \Large [e] = \frac{Coulombs}{electron}\] See, [e] is the charge per electron, so when you multiply n and e together, the "electrons" unit divides out leaving you with coulombs. This is called dimensional analysis, but it's really just a fancy way of saying we're looking to make sure the units work out. --- To answer your question just asked now, yes, the charge on the electron is a constant. All electrons have the exact same charge, that's one of their properties, kind of like how bananas are all yellow. (ok some are green, but whatever you get the idea haha)

OpenStudy (anonymous):

ok, so for number of electrons i'm getting 3.855e^10.

OpenStudy (anonymous):

for the second one, i have the equation f(grav)=(G*m1*m2)/r^2. i was wrong about G. it's not 9.8, it's 6.67*10^-11. take the mass of the two balls (m1 and m2), and the r^2 is technically 7 cm since they started at 1.5 cm apart and ended at 8.5 cm apart. is that correct?

OpenStudy (kainui):

Yes, this looks almost perfect, except that you should not consider their initial positions. When calculating force of mass or charge you should always calculate it from their center of mass. So for the two balls, it will likely be something like this:  So their distance for the r^2 term in the bottom should be the distance between the balls and the radius of both balls to get the true distance from center to center.

OpenStudy (anonymous):

so the r^2 is 8.5 since he said to measure from center of ball to center of ball.

OpenStudy (anonymous):

and i have to convert the grams to kg and the cm to meters cause the G is in m/kg

OpenStudy (anonymous):

2.308 *10-17.

OpenStudy (anonymous):

i'm hoping most of our grades come from showing our work in our lab reports. i've never written one and he's not answering student's questions that have been posted all week on the "blackboard". i emailed the dept chair....it's annoying when a teacher doesn't answer your questions. i understand a day or two, but not a whole week with no response in an 8 week course.....

OpenStudy (kainui):

Oh wow an 8 week course, that's intense. Your method is completely correct, so that's good. It looks like you're done.

OpenStudy (anonymous):

yeah, stupid. the lecture part is 16 weeks and the lab is 8. makes no sense at all. thanks for your help.